OpenStudy (anonymous):

A car moving at 40 km /hr. to be stop by applying brakes in the next 4.0 m if the car wt. 2000kg what average force must be applied on it?

OpenStudy (anonymous):

\[v ^{2} = u ^{2} + 2aS\] so \[0 = (40,000/3600)^{2} -2*<X>*4\] gives X = 1250/81 m/s^2 or decelartion of 1250/81 m/s^2 which gives force as 1250/81 * 2000 = 30864.20 N